# sylinline,color=cyan

### 23 Mar 2026 v1

@@ Abstract.
For a smooth projective variety $X\subseteq\mathbb{P}^{N}$ over an algebraically closed field of char 0, we show that the discriminant locus of a generic projection of $X$ is projectively dual to a general linear section of the dual variety, and deduce a purity statement for the discriminant. Over $\mathbb{C}$, we also show that the fundamental group of the complement of the branch divisor arising from generic projection of a normal hypersurface surjects onto a braid group via braid monodromy.
@@ 1. Introduction
where $l_{0},\ldots,l_{k}$ are $k+1$ general linear forms forming a basis of $V$. The two families are related by a linear dual on the ambient projective space. When $k=1$, up to blowing up the base locus, the two families coincide and are precisely the Lefschetz fibration. When $k\geq 2$, they are different: The fiber of generic projection is a higher codimension linear section. We want to explore the relationship between the two families, particularly where the fibers become singular. This is captured by the notion of discriminant.
@@
Let $f\mathrel{\mathop{\ordinarycolon}}\mathcal{Y}\to B$ be a flat family of projective varieties. Assume that $B$ is smooth. Define the discriminant locus $\Delta\subseteq B$ as the closed subspace consisting of points $b$ where the fiber $\mathcal{Y}_{b}$ is singular.
When $f\mathrel{\mathop{\ordinarycolon}}\mathcal{Y}\to B$ is a proper non-dominant map, we use the convention that $\Delta=f(\mathcal{Y})$. See more discussion in Section 3.
For the hyperplane sections family (1), the discriminant locus is precisely the linear section by $\mathbb{P}V$ of the dual variety $X^{\perp}\cap\mathbb{P}V$ [GKZ08], which is reduced as long as $\mathbb{P}V$ is in general position, and is moreover irreducible if it has positive dimension. When $k=2$, a general plane section $\mathbb{P}^{2}\cap X^{\perp}$ is expected to have only ordinary nodes and cusps, then their numbers can be computed explicitly using intersection theory (cf. [DL81, (2.2), (2.3)], [ROB76]).
@@
How to characterize the discriminant locus $\Delta$ of the generic projection of $X$?
In particular, one may ask about its dimension, irreducibility, degree, and singularities. Here we interpret the discriminant of the rational map (2) as the discriminant of the blowup $Bl_{B}X\to\mathbb{P}V^{\vee}$ of the base locus.
@@ 1.1. Duality of discriminant locus
Let $X\subseteq\mathbb{P}^{N}$ be a smooth projective variety, and $\mathbb{P}V$ is a general $k$-dimensional linear subspace in the dual space $(\mathbb{P}^{N})^{\vee}$. Let $\Delta^{\perp}$ denote the dual variety of the discriminant locus of the generic projection (2). Under the natural identification $(\mathbb{P}V^{\vee})^{\vee}\cong\mathbb{P}V$, one has
$\Delta^{\perp}=\mathbb{P}V\cap X^{\perp}.$
@@
When $\pi$ is not dominant (i.e., $\dim(X)<k$), $\pi$ is a generic immersion, and the discriminant coincides with the projection image $\pi(X)$. In this case, a version of Theorem 1.3 is discussed in [TEV05, Theorem 1.21] and [HOL88, Propsition 6.1]; When $X$ is a hypersurface, Theorem 1.3 is consistent with [PIE78, Corollary 4.3]. We are particularly interested in the case where the generic projection is dominant. We also show that the result still holds when $X$ is normal and $\pi$ is finite and dominant (cf. Theorem 5.7).
@@ 1.2. Purity of discriminant locus
Depending on the codimension of the dual variety $X^{\perp}\subseteq(\mathbb{P}^{N})^{\vee}$, the discriminant locus $\Delta(\pi_{k})$ of the generic projection $\pi_{k}$ satisfies one of the following:
@@
    $\Delta(\pi_{k})$ is an irreducible hypersurface, if $\textup{codim}(X^{\perp})<k$;
@@
    $\Delta(\pi_{k})$ is a finite union of hyperplanes, if $\textup{codim}(X^{\perp})=k$;
@@
    $\Delta(\pi_{k})$ is empty, if $\textup{codim}(X^{\perp})>k$.
@@
A generalization of Theorem 1.3 and 1.4 still hold without smoothness assumption of $X$, but the claims only apply to a closed subspace of the discriminant captured by the smooth locus of $X$ (see Definition 3.1, Theorem 4.1 and 5.3).
1.3. Degree of discriminant
A projective variety $X$ and its dual $X^{\perp}$ is related by a natural correspondence:
$X\xleftarrow{\pi_{1}}N_{\mathbb{P}^{N}}^{\perp}\xrightarrow{\pi_{2}}X^{\perp},$
where $N_{\mathbb{P}^{N}}^{\perp}$ is the conormal variety of $X$ and lives in $\mathbb{P}^{N}\times(\mathbb{P}^{N})^{\vee}$. For a general $\mathbb{P}V\subseteq(\mathbb{P}^{N})^{\vee}$, the variety $\pi_{1}\big(\pi_{2}^{-1}(\mathbb{P}V\cap X^{\perp})\big)$ is called the polar locus in the literature, whose degree is computed in [KLE86]. We show that the discriminant locus $\Delta$ of a generic projection is the projection image of the polar locus (which is called smooth critical locus in our Definition 3.1). We show that this map is birational, hence they have the same degree and we obtain the following
Proposition 1.7.
The discriminant locus $\Delta_{k}$ of the generic projection (3) has degree
$\deg(\Delta_{k})=\int_{N^{\perp}_{\mathbb{P}^{N}}(X)}\pi_{1}^{*}H^{k-1}\cdot\pi_{2}^{*}(H^{\prime})^{N-k},$
where $H=c_{1}\mathcal{O}_{\mathbb{P}^{N}}(1)$ and $H^{\prime}=c_{1}\mathcal{O}_{(\mathbb{P}^{N})^{\vee}}(1)$ are hyperplane classes.
We use Theorem 1.3 to deduce some classical results on the degree of polar loci. See Section 8.
@@ 1.3. Degree of discriminant
For varieties over $\mathbb{C}$, fundamental groups of complements of hypersurfaces are closely related to the geometry of singular hypersurfaces and have been studied for a long time [ZAR29, KAM33, MT88, DL81, DIM92]. In some of the interesting examples, it’s computed that the fundamental groups of hypersurface complements are isomorphic to braid groups, $B_{m}$, which are roughly speaking the group of trajectories of a finite collection of points moving on the sphere $\mathbb{S}^{2}\cong\mathbb{P}^{1}$ (see Section 9).
We extended the investigation to generic finite projection of a normal hypersurface, and relate to special dual varieties.
@@
Let $Y\subseteq\mathbb{P}^{N+1}$ be a normal irreducible hypersurface of degree $m$. Let $Y\to\mathbb{P}^{N}$ be the generic linear projection (note this is a finite branched cover). Let $\Delta$ be the branch divisor.
  • Then $\deg(\Delta)=m(m-1)$, and there exists a surjective braid monodromy homomorphism
@@
  • When $Y$ is the dual variety of a smooth projective variety $\mathcal{X}$, then the branch divisor $\Delta$ is identified with $X^{\perp}$, where $X=H\cap\mathcal{X}$ is a general hyperplane section. In particular, there is a surjection
@@
Remark 1.9.
Varieties $\mathcal{X}$ with normal dual varieties are studied by Zak [ZAK91]. It is shown that such condition guarantees that $\mathcal{X}$ is non-extendable, i.e., not a smooth hyperplane section of one-dimensional higher, if it is not a quadric hypersurface. Examples of such $\mathcal{X}$ include degree two Veronese embedding of $\mathbb{P}^{n}$ (see $n=2$ below) and Severi varieties [ZAK93].
Example 1.10.
Let $\mathcal{X}=v_{2}(\mathbb{P}^{2})$ be the 2nd Veronese embedding of $\mathbb{P}^{2}$. Then $\mathcal{X}^{\perp}$ is a normal cubic hypersurface (see Section 7.1). A general hyperplane section $X=\mathcal{X}\cap H$ is the rational normal quartic, and $X^{\perp}$ is the hypersurface of degree $6$ parameterizing the set of unordered $4$ points on $\mathbb{P}^{1}$ with at least two points colliding, using the identification $\mathbb{P}^{4}\cong\textup{Sym}^{4}(\mathbb{P}^{1})$. A general plane section $X^{\perp}\cap\mathbb{P}^{2}$ is a rational curve with $4$ nodes and $6$ cusps. This is a maximal cuspidal rational curve studied by Zariski [ZAR36]. It is shown that the fundamental group of the complement $\pi_{1}(\mathbb{P}^{2}\setminus X^{\perp},*)$ is the braid group $B_{4}$. Hence the braid monodromy representation from Theorem 1.8 gives surjection $B_{4}\twoheadrightarrow B_{3}$.
@@
In Section 3, we introduce the notion of smooth discriminant, and discuss its relationship with the classical discriminant 1.1. We recall Zariski–Nagata theorem for finite branched cover.
@@
Section 6, 7 and 8 are applications of the main theorems. In Section 6, we prove Corollary 1.5 and discuss failure in some singular setting. In Section 7, we demonstrate the duality theorem 1.3 in surfaces and threefolds, and compute singularities of discriminant curves. In Section 8, we discuss the relationship to the polar locus. We find the degree of the smooth discriminant of a generic projection and prove Proposition 1.7.
@@
Acknowledgment. We would like to thank Anatoly Libgober, Christian Schnell, and John Sheridan for helpful conversations and correspondence. We are grateful to Thomas Dedieu for pointing out the distinction between our definition of smooth discriminant (Definition 3.1) and the definition via Fitting ideals in the literature, for carefully reading an earlier draft, and for many helpful comments on the revision. We gratefully acknowledge the support and hospitality of the Simons Center for Geometry and Physics at Stony Brook University during the 2nd Simons Math Summer Workshop: Moduli in July 2024, where a project motivating this work began. S.-Y. Liu is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - GZ 2047/1, Projekt-ID 390685813.
@@ 3. Discriminant of an algebraic map
The Definition 1.1 of discriminant of a proper flat family $\pi\mathrel{\mathop{\ordinarycolon}}\mathcal{Y}\to B$ uses $d$-th Fitting ideal, where $d=\dim(\mathcal{Y})-\dim(B)$ is the relative dimension [TEI77] [STA18, 0C3K]. In this section, we use $\Delta_{Fitt}(\pi)$ to denote this discriminant. $\Delta_{Fitt}(\pi)$ has a scheme structure, but in this paper, we only focus on the set-theoretical properties of this discriminant and treat it as a variety, with the reduced scheme structure.
@@ 3.1. Smooth discriminant
    is not surjective, and $\Delta(\pi)$ is the classical discriminant (cf. Definition 1.1).
@@
    Note that when $\pi$ is not dominant, $\Delta(\pi)=\pi(\mathcal{Y})$ is the entire image, and this is different from the conventions in the literature, e.g., [SHI06, Definition 2.15].
@@
Let $\pi\mathrel{\mathop{\ordinarycolon}}\mathcal{Y}\to B$ be a proper morphism and assume that $B$ is smooth. Then set-theoretically, the discriminant $\Delta_{Fitt}(\pi)$ is the union of the smooth discriminant and the image of the singularities of the total space:
@@ 3.2. Finite branched cover
Let $\pi\mathrel{\mathop{\ordinarycolon}}\mathcal{Y}\to B$ be a finite dominant morphism of varieties. Assume $B$ is smooth and $\mathcal{Y}$ is normal. Let $U\subset B$ be the maximal open subset such that $\pi^{-1}(U)\to U$ is étale, and set $\Delta_{Fitt}(\pi)\mathrel{\mathop{\ordinarycolon}}=B\setminus U$. Then $\Delta_{Fitt}(\pi)$ is either empty or has pure codimension one in $B$.
@@
For a dominant proper morphism $\pi\mathrel{\mathop{\ordinarycolon}}\mathcal{Y}\to B$, we call its discriminant clean if $\pi(Sing(\mathcal{Y}))$ is contained in the smooth discriminant, in other words, there is a set theoretical identity
@@
be the generic projection (note this is a finite branched cover). Then the discriminant $\Delta(\pi)$ is clean.
@@ 4. Duality of smooth discriminant
In this section, we will prove Theorems 1.3. In fact, we will prove a general statement (Theorem 4.1) for projective duality of the smooth discriminant of a generic projection of an irreducible projective variety $X\subseteq\mathbb{P}^{N}$ over an algebraically closed field with char 0. $X$ is not necessarily smooth.
@@ 4.1. The setup
Throughout this section, we use an opposite convention from the introduction—We take a linear section of $X$ and a generic projection of the dual variety $X^{\perp}$ to a lower-dimensional subspace. (Note that by classical projective duality, we can switch the roles of $X$ and $X^{\perp}$.)
@@
Note that Theorem 1.3 follows from the above, together with classical projective duality 2.5 and Lemma 3.3. To prove Theorem 4.1, the key is to describe the smooth critical locus of the generic projection $\pi$ using an incidence relation (cf. Proposition 4.7). Then the proof follows from classical projective duality and the generality lemma for the choice of $\mathbb{P}V$.

### 19 Apr 2026 v2

@@ Abstract.
For a smooth projective variety $X\subseteq\mathbb{P}^{N}$ over an algebraically closed field of char 0, we show that the discriminant locus of a generic projection of $X$ is projectively dual to a general linear section of the dual variety, and deduce a purity statement for the discriminant. Over $\mathbb{C}$, we also show that the fundamental group of the complement of the branch divisor of a generic projection of a normal hypersurface surjects onto a braid group via braid monodromy.
@@ 1. Introduction
where $l_{0},\ldots,l_{k}$ are $k+1$ general linear forms forming a basis of $V$. The base locus is $X\cap\mathbb{P}V^{\perp}$. The two families are related by a linear dual on the ambient projective space. When $k=1$, up to blowing up the base locus, the two families coincide and are precisely the Lefschetz fibration. When $k\geq 2$, they are different: The fiber of generic projection is a higher codimension linear section. We want to explore the relationship between the two families, particularly where the fibers become singular. This is captured by the notion of discriminant.
@@
Let $f\mathrel{\mathop{\ordinarycolon}}\mathcal{Y}\to B$ be a proper dominant morphism. Assume that $B$ is smooth. Define the discriminant locus $\Delta\subseteq B$ as the closed subspace consisting of points $b$ where the fiber $f^{-1}(b)=\mathcal{Y}_{b}$ is singular. In the flat setting, this agrees set-theoretically with the usual definition via Fitting ideal. When $f\mathrel{\mathop{\ordinarycolon}}\mathcal{Y}\to B$ is a proper non-dominant map, we use the convention that $\Delta=f(\mathcal{Y})$. See more discussion in Section 3.
For the hyperplane sections family (1), the discriminant locus is precisely the linear section by $\mathbb{P}V$ of the dual variety $\mathbb{P}V\cap X^{\perp}$ [GKZ08], which is reduced as long as $\mathbb{P}V$ is in general position, and is moreover irreducible if it has positive dimension. When $k=2$, a general plane section $\mathbb{P}^{2}\cap X^{\perp}$ is expected to have only ordinary nodes and cusps, then their numbers can be computed explicitly using intersection theory (cf. [DL81, (2.2), (2.3)], [ROB76]).
@@
How can one characterize the discriminant locus $\Delta$ of the generic projection of $X$?
In particular, one may ask about its dimension, irreducibility, degree, and singularities. Here we interpret the discriminant of the rational map (2) as the discriminant of the regular morphism $\tilde{X}\to\mathbb{P}V^{\vee}$ obtained by blowing up the base locus.
Our approach is related to the classical theory of polar varieties and generic projections, especially the work of Piene [PIE78], Kleiman [KLE86], Lê–Teissier [TT88], and Flores–Teissier [FT18]. In particular, the relation between critical loci of projections and polar loci already appears in that literature in various forms. The new point here is to reinterpret this relation through projective duality, identifying the discriminant of a generic projection with the projective dual of a general linear section of $X^{\perp}$, and to deduce from this a purity statement and further applications.
@@ 1.1. Duality of discriminant locus
Let $X\subseteq\mathbb{P}^{N}$ be a smooth projective variety, and $\mathbb{P}V$ is a general $k$-dimensional linear subspace in the dual space $(\mathbb{P}^{N})^{\vee}$. Then under the natural identification $(\mathbb{P}V^{\vee})^{\vee}\cong\mathbb{P}V$, one has
$\Delta=(\mathbb{P}V\cap X^{\perp})^{\perp}.$
@@
Note that in the extreme case when $k=N$, the statement reduces to the classical projective duality. Thus Theorem 1.3 can be thought of as a relative version of projective duality.
When $\pi$ is not dominant, i.e. when $\dim(X)<k$, the map $\pi$ is a generic immersion onto its image, and the discriminant coincides with the projection image $\pi(X)$. In this case, a version of Theorem 1.3 appears in [TEV05, Theorem 1.21] and [HOL88, Proposition 6.1]. We are mainly interested in the dominant case.
@@ 1.2. Purity of discriminant locus
Let $c$ be the codimension of the dual variety $X^{\perp}$ in $(\mathbb{P}^{N})^{\vee}$, then
@@
    if $c<k$, $\Delta(\pi_{k})$ is an irreducible hypersurface;
@@
    if $c=k$, $\Delta(\pi_{k})$ is a finite union of hyperplanes;
@@
    if $c>k$, $\Delta(\pi_{k})$ is empty.
@@
We also show that both Theorem 1.3 and 1.4 remain valid when $X$ is normal and $\pi$ is finite and dominant; see Theorem 5.7. When $X$ is singular in general, the claims about duality and purity still apply to a closed subspace of the discriminant captured by the smooth locus of $X$ (see Definition 3.1, Theorem 4.1 and 5.3).
@@
For varieties over $\mathbb{C}$, fundamental groups of complements of hypersurfaces are closely related to the geometry of singular hypersurfaces and have been studied for a long time [ZAR29, KAM33, MT88, DL81, MOI81, DIM92]. In some of the interesting examples, it’s computed that the fundamental groups of hypersurface complements are isomorphic to braid groups, $B_{m}$, which are roughly speaking the group of trajectories of a finite collection of points moving on the sphere $\mathbb{S}^{2}\cong\mathbb{P}^{1}$ (see Section 9).
We extend this perspective to generic finite projections of normal hypersurfaces and relate the resulting branch divisors to projective dual varieties of hyperplane sections.
@@
Let $Y\subseteq\mathbb{P}^{N+1}$ be a normal hypersurface of degree $m$, and let $\pi\mathrel{\mathop{\ordinarycolon}}Y\to\mathbb{P}^{N}$ be a general linear projection with branch divisor $\Delta$. Then:
  • $\Delta$ is a hypersurface of degree $\deg(\Delta)=m(m-1)$. If $N>1$, then $\Delta$ is irreducible;
  • let $\mathcal{X}=Y^{\perp}\subseteq(\mathbb{P}^{N+1})^{\vee}$, and $X=H\cap\mathcal{X}$ is a general hyperplane section, then $\Delta\cong X^{\perp}$;
  • the braid monodromy induces a surjection
@@
Remark 1.8.
Moishezon [MOI81] proved that when $Y$ is smooth and more generally if $Sing(Y)$ has codimension at least 3, then the braid monodromy representation (4) is actually an isomorphism. This fails in general for normal $Y$; see Example 9.8.
Corollary 1.9.
Let $\mathcal{X}$ be a smooth projective variety whose dual variety is a normal hypersurface of degree $m$. Let $X=H\cap\mathcal{X}$ be a general hyperplane section. Use the identification $\mathbb{P}^{N}\cong H^{\vee}$, then there is a surjection
@@
Smooth varieties with normal dual were studied by Zak [ZAK91]. The normality condition on the dual forces $\mathcal{X}$ to be non-extendable unless $\mathcal{X}$ is a quadric hypersurface. Examples of such $\mathcal{X}$ include determinantal varieties, such as the second Veronese embedding $v_{2}(\mathbb{P}^{n})$, the Severi varieties [ZAK93], and the Segre embeddings of $\mathbb{P}^{n}\times\mathbb{P}^{n}$ of bidegree $(1,1)$.
@@
In Section 3, we introduce the notion of smooth discriminant and compare it with the discriminant of Definition 1.1. We also recall the Zariski–Nagata purity theorem for finite branched covers.
@@
Section 6, 7 and 8 are applications of the main theorems. In Section 6, we prove Corollary 1.5 and discuss failure in some singular setting. In Section 7, we demonstrate Theorem 1.3 in surfaces and threefolds, and compute singularities of discriminant curves. In Section 8, we discuss the relationship to the polar locus and deduce some classical results on the degree of polar loci.
@@
Acknowledgment. We would like to thank Anatoly Libgober, Christian Schnell, John Sheridan, and Bernard Teissier for helpful conversations and correspondence. We are grateful to Thomas Dedieu for pointing out the distinction between our definition of smooth discriminant (Definition 3.1) and the definition via Fitting ideals in the literature, for carefully reading an earlier draft, and for many helpful comments on the revision.
We gratefully acknowledge the support and hospitality of the Simons Center for Geometry and Physics at Stony Brook University during the 2nd Simons Math Summer Workshop: Moduli in July 2024, where a project motivating this work began. S.-Y. Liu is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - GZ 2047/1, Projekt-ID 390685813.
@@ 3. Discriminant of an algebraic map
For a proper flat family $\pi\mathrel{\mathop{\ordinarycolon}}\mathcal{Y}\to B$, the discriminant of Definition 1.1 agree set-theoretically with the discriminant defined via the $d$-th Fitting ideal of $\Omega_{\mathcal{Y}/B}$, where $d=\dim(\mathcal{Y})-\dim(B)$ is the relative dimension. See [TEI77] [STA18, 0C3K]. In this section, we use $\Delta_{Fitt}(\pi)$ to denote this discriminant. In this paper, we only focus on the set-theoretical properties of this discriminant and treat $\Delta_{Fitt}(\pi)$ as a variety.
@@ 3.1. Smooth discriminant
    is not surjective. If $\pi$ is moreover flat, then $\Delta(\pi)=\Delta_{Fitt}(\pi)$ is the classical (Fitting) discriminant (cf. Definition 1.1).
@@
    When $\pi$ is not dominant, $\Delta(\pi)=\pi(\mathcal{Y})$ is the image of $\mathcal{Y}$. This is different from the conventions in the literature, e.g., [SHI06, Definition 2.15].
  • (3)
    A similar notion of smooth critical locus $\mathcal{C}(\pi)$ has also been studied by Flores and Teissier in the affine complex analytic setting [FT18].
@@
Let $\pi\mathrel{\mathop{\ordinarycolon}}\mathcal{Y}\to B$ be a proper flat morphism and assume that $B$ is smooth. Then set-theoretically, the discriminant $\Delta_{Fitt}(\pi)$ is the union of the smooth discriminant and the image of the singularities of the total space:
@@ 3.2. Finite branched cover
Let $\pi\mathrel{\mathop{\ordinarycolon}}\mathcal{Y}\to B$ be a finite surjective morphism where $B$ is a smooth variety and $\mathcal{Y}$ is a normal variety. Let $U\subseteq B$ be the maximal open subset such that the restricted morphism $\pi^{-1}(U)\to U$ is étale. Then the branch locus $B\setminus U$ agrees set-theoretically with $\Delta_{Fitt}(\pi)$. Moreover, $\Delta_{Fitt}(\pi)$ is either empty or is a divisor of pure codimension one in $B$.
@@
For a dominant proper morphism $\pi\mathrel{\mathop{\ordinarycolon}}\mathcal{Y}\to B$, we call its discriminant clean if $\pi(Sing(\mathcal{Y}))$ is contained in the smooth discriminant, in other words, if set-theoretically one has
@@
be the generic projection (note this is a finite dominant morphism). Then the discriminant $\Delta(\pi)$ is clean.
@@ 4. Duality of smooth discriminant
In this section, we prove Theorems 1.3. In fact, we will prove a general statement (Theorem 4.1) for projective duality of the smooth discriminant of a generic projection of an irreducible projective variety $X\subseteq\mathbb{P}^{N}$ over an algebraically closed field with char 0. $X$ is not necessarily smooth.
@@ 4.1. The setup
Throughout this section, we use the convention opposite to that of the introduction: We take a linear section of $X$ and a generic projection of the dual variety $X^{\perp}$ to a lower-dimensional linear subspace. (Note that by classical projective duality, we can switch the roles of $X$ and $X^{\perp}$.)
@@
Note that Theorem 1.3 follows from the above, together with classical projective duality 2.5. To prove Theorem 4.1, the key is to describe the smooth critical locus of the generic projection $\pi$ using an incidence relation (cf. Proposition 4.7). Then the proof follows from classical projective duality and the generality lemma for the choice of $\mathbb{P}V$.